# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 4.1.9. Data Split Ratio
Optimal data splitting in the context of AI involves dividing a dataset into training, validation, and testing subsets to maximize the performance and generalization of the AI model. The goal is to allocate the appropriate proportions of data for effective model training, hyperparameter tuning, and unbiased evaluation.
Example: In the context of sentiment analysis, consider a dataset of 10,000 customer reviews to build an AI model for predicting whether a review is positive or negative. To ensure the best performance and generalization of the model, the dataset could be divided into three subsets. The training set comprises 7,000 instances, accounting for 70% of the data, allowing the model to learn and capture underlying patterns and relationships in the reviews. For fine-tuning and hyperparameter selection, 1,500 instances are allocated (15% of the data) to the validation set. This smaller subset helps optimize the model’s performance during training. Lastly, 1,500 instances (15% of the data) are set aside for the testing set. This subset consists of unseen reviews and objectively evaluates the model’s performance, ensuring its ability to generalize effectively to new and real-world instances. By splitting the dataset in this manner, one can ensure that the AI model is trained on diverse and representative data, fine-tuned for optimal performance, and tested on unseen instances, enabling a robust sentiment analysis system.
Metrics in Literature: Affendras et al. (Afendras and Markatou, 2019) suggests that irrespective of data distribution or analytic task, the optimal training sample size in cross-validation is identified as half of the total sample size. Building on this, Joseph (Joseph, 2022) examines the ideal data splitting ratio for training and validation sets in linear regression models. The authors propose a ratio of $\sqrt{p}:1$, where $p$ is the number of parameters required to estimate a well-fitting linear regression model accurately. The authors also present a practical strategy for determining $p$ using variable selection methods. It suggests that this approach can be useful for regression and classification tasks.
Summary: The data split ratio, crucial for optimizing AI model performance, entails dividing a dataset into training, validation, and testing subsets. Affendras et al. suggest a guideline, proposing the training set to be half of the total dataset. With metrics like the $\sqrt{p}:1$ ratio guides to ideal splits in linear regression models.
@@ 4.1.10. Data Point Impact
It refers to the measure of the influence or significance of individual data points within a dataset in the context of AI. It quantifies the extent to which each data point contributes to an AI model or system’s overall performance, accuracy, or behavior.
Example: Consider a dataset containing medical records of patients. Each data point represents a patient’s age, medical history, symptoms, and diagnostic outcomes. By analyzing the data point impact, we can determine which specific patient records have a higher influence on the predictions or outcomes of an AI model built for disease diagnosis.
Consider a machine learning model trained to predict the likelihood of a patient having a specific medical condition based on the dataset. By calculating the data point impact, we can identify the patient records with the most significant effect on the model’s predictions. These influential data points could correspond to patients with rare symptoms or unique characteristics that strongly impact the model’s decision-making process.
Metrics in Literature: Ghorbani et al. (Ghorbani and Zou, 2019) introduced the Data Shapley metric, based on the Shapley value from cooperative game theory, to assess the value of individual data points in supervised machine learning. The metric measures the contribution of each data point to the model’s predictions, revealing its importance in model training. Various techniques, such as Monte Carlo and gradient-based methods, estimate a data point’s value by considering its combinations with different subsets of the training data. Similarly, Wang et al. (Wang and Jia, 2023) propose the Banzhaf value, a metric to assess data point value in the presence of noisy model performance scores. The authors investigate the robustness of data valuation in stochastic gradient descent, where randomness can lead to inconsistent data value rankings.
Summary: Data point impact refers to the measure of the influence or significance of individual data points within a dataset. One commonly used metric to evaluate data point impact is the Data Shapley metric. It calculates the value of individual data points based on their contribution to the model’s predictions. The Banzhaf value is another metric that addresses the challenge of noisy model performance scores and provides robust differentiation of data quality in stochastic machine learning algorithms.
@@ 4.1.11. Correctness
In the context of data values for AI, it refers to the degree of accuracy and fidelity in representing real-world information. It measures how closely the recorded data values align with the actual values they intend to represent. The goal is to minimize discrepancies, errors, or deviations between the recorded data and the ground truth.
Example: Consider a dataset containing temperature measurements from weather stations. The correctness of data values would involve ensuring that each recorded temperature value accurately reflects the actual temperature at the corresponding location and time. Variations, inconsistencies, or inaccuracies in the recorded values compared to the actual temperature values would indicate a lack of correctness in the dataset.
Metrics in Literature: Pipino et al’s(Pipino et al., 2002) correctness metric quantifies data accuracy by calculating the complement of the error ratio. It emphasizes clear criteria, like precision levels, and acknowledges contextual variations in error tolerability for a systematic evaluation of data correctness. Similarly, Kaiser et al. (Kaiser et al., 1970) involves comparing attribute values in the dataset ($w_{I}$) with their corresponding values in the real world ($w_{R}$). A domain-specific distance function, denoted as $d(w_{I},w_{R})$, quantifies the difference between these attribute values. The metric’s objective is to ensure normalization within the interval [0, 1] without using a quotient.
Summary: Correctness, in the context of data values for AI, refers to the accuracy and fidelity of the recorded data in representing real-world information. They involve calculations related to error ratios, precision criteria, contextual error tolerability, and comparisons between dataset attribute values and their real-world counterparts.
@@ 4.1.12. Timeliness
The timeliness of data refers to the recency and relevance of the data in relation to the current state of the phenomenon or domain it represents. It measures how closely the data captures the most recent information available at the time of analysis or model training, ensuring that the data is up-to-date and reflects the present conditions.
@@
Example: Consider an AI system that predicts product demand for an e-commerce company. Timeliness of the data used for training the model would involve using the most recent sales data, customer preferences, and market trends. If the dataset contains sales data from several months ago, it may not accurately capture the current demand patterns and consumer behavior. By ensuring timely data, such as incorporating daily or weekly sales updates, the AI model can better adapt to the changing market dynamics and provide more accurate demand predictions.
@@
Metrics in Literature: Two studies, one by Kaiser et al. (Kaiser et al., 1970) and the other by Heinrich et al. (Heinrich and Klier, 2015), introduce data readiness metrics for evaluating the timeliness of attribute values. Both metrics use probability-based approaches to assess the freshness and relevance of data. Kaiser et al.’s metric employs an exponential distribution model to calculate attribute decline rates, indicating the average proportion of outdated attribute values within a specified time frame. It quantifies attribute age based on data quality assessment time and data acquisition time, offering an automated and interpretable measure of timeliness. In contrast, Heinrich et al. propose a probability-based currency metric (PBCM) that assesses data item timeliness using a set of probabilities. These probabilities are derived from diverse data sources and methods, including expert assessments, historical data analysis, and machine learning algorithms. The PBCM combines these probabilities using a weighted average formula to generate an overall score, with weights determined by context and importance. While both metrics share a foundation in probability theory, Kaiser et al.’s metric focuses on attribute-level timeliness. In contrast, Heinrich et al.’s PBCM assesses data item timeliness, offering flexibility for various data types and contexts.
@@
Blake et al. (Blake and Mangiameli, 2011) propose a method to assess data timeliness using a classification model. They introduce a metric called $T$, which measures the impact of introducing new and more current data into the dataset. To evaluate data volatility and timeliness, the authors replace a percentage of old instances with new ones in the training data while assuming a fixed currency. The $T$ metric is computed based on the total quantity of records in the training and test data and the number of replacement records introduced for reclassification. This metric provides valuable insights into the timeliness of data and its impact on the performance of the classification model.
@@
Summary: The timeliness metric for data readiness in AI assesses the recency and relevance of the data about the current state of the phenomenon or domain it represents. Kaiser et al. and Heinrich et al. introduced metrics for assessing data timeliness. Both use probability-based approaches, with Kaiser et al. focusing on attribute-level timeliness and Heinrich et al. assessing data item timeliness. Another method by Blake et al. evaluates data timeliness through a classification model by assessing data replacement’s impact on model performance.
@@ 4.1.13. Privacy Leakage
Data privacy in the context of AI refers to protecting and preserving sensitive information contained within datasets, particularly concerning the risk of unauthorized disclosure or inference of private details. One notable technique used to assess privacy is Membership Inference Attacks (MIA).
@@
Membership Inference Attacks involve determining whether a specific data record was included in the training dataset used to build an AI model. By exploiting patterns and characteristics of the model’s outputs, one can infer whether or not a particular data point was part of the training set. This raises concerns about the privacy of individual data records and the potential for unauthorized access to such information.
@@
Example: Consider a healthcare dataset used to train a machine-learning model for disease diagnosis. The dataset contains sensitive medical information about patients, including their symptoms, test results, and diagnoses. Suppose one can successfully perform a membership inference attack on this model. In that case, one may determine whether a specific patient’s data was used during training, posing a significant privacy risk as it could reveal a patient’s medical condition or other confidential information.
@@
Metrics in Literature: The focus of Vatsalan et al.’s (Vatsalan et al., 2022) work centers on mitigating re-identification risks in released datasets within the education sector. In contrast to existing approaches that often presume prior knowledge, the proposed method employs a Markov Model to quantify re-identification risks by leveraging all available information in the datasets, including event-level details associating multiple records with the same individual and exploring correlations between attributes.
@@
In the broader context of privacy metrics in literature, the study positions itself alongside notable works such as SHAP R introduced by Duddu et al. (Duddu et al., 2022) and Song et al.’s (Song and Mittal, 2021) privacy risk metric. SHAP R quantifies the susceptibility of individual training data records to membership inference attacks by calculating Shapley values, emphasizing the influence of specific data points on model predictions. In contrast, Song et al.’s metric assesses the likelihood of a data record being present in a model’s training dataset, focusing on evaluating the privacy risk from an adversarial perspective. Both metrics aim to address privacy concerns by assessing the privacy risks associated with data records, with their applicability tied to a specific AI model.
@@
In another study, Carlini et al. (Carlini et al., 2022) introduced the Attack Success Rate (ASR) as a metric for privacy leakage. ASR measures the success of an attack in predicting if a specific example is part of the training dataset. It is calculated by training a model, performing an attack, and evaluating the attack’s success in correctly predicting membership. However, ASR can only be measured after training a model and conducting an attack. Alongside these contributions, entropy-based metrics, demonstrated by Bezzi (Bezzi, 2007), Longpr ’e et al. (Longpré et al., 2017), and Sevgi et al. (Arca and Hewett, 2020) offer valuable insights into dataset anonymity by leveraging entropy to measure unpredictability and disorder in released data.
@@
Summary: The privacy leakage metric in AI aims to safeguard sensitive information from unauthorized disclosure. Vatsalan et al. employ a Markov Model to address re-identification risks, while SHAP R quantifies membership privacy risk using Shapley values. Attack Success Rate (ASR) gauges privacy leakage in membership inference attacks. Entropy-based metrics are also explored for assessing dataset anonymity.
@@ 4.1.14. Sample Size
In the context of AI, it refers to the number of data points or instances selected from a population to be included in a dataset for analysis or training. It represents the subset of data used to make inferences or predictions.
@@
Example: Consider a medical research study to predict a particular disease’s likelihood based on various patient characteristics. The researchers collected data from 500 patients, including their age, gender, medical history, and test results. In this case, the sample size of the dataset would be 500, indicating that the analysis or model development is based on these 500 patient records.
The sample size is essential in AI as it can affect the statistical power, generalizability, and reliability of the results or models. A larger sample size generally provides more representative and reliable insights, while a smaller sample size may introduce higher uncertainty or limitations in the analysis.
@@
Metrics in Literature: In their study, Alwosheel et al. (Alwosheel et al., 2018) investigate the sample size requirements for accurate decision-making analysis using Artificial Neural Networks (ANNs). They propose a novel guideline called “factor 50,” suggesting that the ideal dataset size for an ANN should be the number of adjustable parameters in the model multiplied by 50, which is more conservative compared to the widely used “factor 10” rule-of-thumb found in the literature (Haykin, 2009). However, determining the appropriate sample size for ANN-based decision-making is complex and depends on the model’s complexity, which is difficult to predict. To address this, the authors propose three approaches: evaluating ex-post if the training sample size was sufficient, using prior studies or literature to estimate the optimal number of neurons, or leveraging existing literature to estimate the expected number of neurons required for the analysis.
@@
Summary: In AI, sample size denotes the quantity of data points chosen from a population for analysis or model training. Alwosheel et al. propose the “factor 50” guideline for determining sample size in decision-making with Artificial Neural Networks (ANNs), contrasting with the commonly used “factor 10”.
4.1.15. FAIRness Score

### v2

@@ 3.1.9. Data Split Ratio
Optimal data splitting in AI involves dividing a dataset into training, validation, and testing subsets to maximize the performance and generalization of the AI model. This metric aims to allocate the appropriate proportions of data for effective model training, hyperparameter tuning, and unbiased evaluation.
Example: Datasets are typically split with a ratio of $60$/$20$/$20$ for training, validation, and testing. Split ratios of $70$/$15$/$15$, $80$/$10$/$10$ are also common. By splitting the dataset in this manner, we can ensure that the AI model is trained on diverse and representative data, fine-tuned for optimal performance, and tested on unseen instances, enabling a robust sentiment analysis system.
Metrics in Literature: Afendras and Markatou (2019) suggests that irrespective of data distribution or analytic task, the optimal training sample size in cross-validation is identified as half of the total sample size. Similarly, Joseph (2022) examines the ideal data splitting ratio for training and validation sets in linear regression models. The authors propose a ratio of $\sqrt{p}:1$, where $p$ is the number of parameters required to estimate a well-fitting linear regression model. The authors also present a strategy for determining $p$ using variable selection methods. It suggests that this approach can be helpful in regression and classification tasks.
Impact on AI: An optimal split ensures that the model is trained on a sufficient amount of data to learn effectively while being validated and tested on separate, representative subsets to evaluate its generalization capabilities. An inappropriate split ratio can lead to overfitting or underfitting. Additionally, ensuring that the split maintains the statistical distribution of the data is crucial to avoid biases.
Summary: The data split ratio, which involves dividing a dataset into training, validation, and testing subsets, is crucial for optimizing AI model performance. Affendras et al. suggest a guideline, proposing the training set to be half of the total dataset. With metrics like the $\sqrt{p}:1$ ratio guide to ideal splits in linear regression models.
@@ 3.1.10. Data Point Impact
It refers to the measure of the influence or significance of individual data points within a dataset. It quantifies the extent to which each data point contributes to an AI model or system’s overall performance, accuracy, or behavior.
Example: Consider a patient medical record dataset with the patient’s age, medical history, symptoms, and diagnostic outcomes. By analyzing the impact of data points, we can determine which specific patient records have a higher influence on the outcomes of an AI model built for disease diagnosis.
Metrics in Literature: Ghorbani and Zou (2019) introduced the Data Shapley metric, based on the Shapley value from cooperative game theory. It assesses the impact of individual data points in supervised machine learning. The metric measures the contribution of each data point to the model’s predictions, revealing its importance in model training. Various techniques, such as Monte Carlo and gradient-based methods, estimate a data point’s impact by considering its combinations with different subsets of the training data. Similarly, Wang and Jia (2023) propose the Banzhaf value, a metric to assess data point value in the presence of noisy model performance scores. The authors investigate the robustness of data valuation in stochastic gradient descent, where randomness can lead to inconsistent value rankings.
In addition to these, several other methods have been developed to measure data importance, including the Leave-One-Out (LOO) (Cook and Weisberg, 1982) evaluation, which assesses model performance changes when individual data points are removed, influence functions (Koh and Liang, 2017) estimate a data point’s effect based on loss function gradients, and k-nearest neighbors (KNN) approaches analyze proximity to decision boundaries. Core-set selection (Bachem et al., 2017) identifies impactful subsets that perform similarly to the full dataset. Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016) provide insights by approximating complex models with interpretable ones around specific predictions.
Impact on AI: Evaluating the influence of specific data points allows for better understanding and optimization of the model, ensuring that the most relevant and informative data is used while minimizing noise and irrelevant information. This process can improve the accuracy, efficiency, and generalization capabilities of AI systems, leading to more reliable predictions and insights. Additionally, understanding the impact of data points helps identify and mitigate biases, ensuring fair and equitable AI outcomes.
Summary: Data point impact refers to the measure of the influence or significance of individual data points within a dataset. In addition to feature importance metrics, such as Shapley and Banzhaf values and LIME, influence functions and ablation (removing data points) approaches like LOO are used to measure data point impact.
@@ 3.1.11. Correctness
In terms of data values for AI, correctness refers to the degree of accuracy and fidelity in representing the information of the system being analyzed. It measures how closely the recorded data values align with the actual values they are supposed to represent. The goal of the metric is to minimize discrepancies between the recorded data and the ground truth.
Example: Consider a dataset containing temperature measurements from weather stations. The correctness of data values would involve ensuring that each recorded temperature value accurately reflects the actual temperature at the corresponding location and time. Inaccuracies in the recorded values compared to the actual temperature values would indicate a lack of correctness in the dataset.
Metrics in Literature: Pipino et al’s(Pipino et al., 2002) correctness metric quantifies data accuracy by calculating the complement of the error ratio. It focuses on clear criteria, like precision levels, and recognizes the contextual variations in error tolerance. This ensures a systematic evaluation of data correctness. Similarly, Kaiser et al. (1970) involves comparing attribute values in the dataset ($w_{I}$) with their corresponding values in the real world ($w_{R}$). A domain-specific distance function, denoted as $d(w_{I},w_{R})$, quantifies the difference between these attribute values. The objective of the metric is to ensure normalization within the interval $[0,1]$ without using a quotient.
Impact on AI: Correct and accurate data ensures that AI models can learn true patterns and make precise predictions to reduce the likelihood of errors and biases. Inaccurate data can lead to flawed outcomes, such as false positives or negatives, undermining the effectiveness of the AI application by spoiling user trust. It also enhances the generalization and robustness of models, enabling them to handle diverse inputs and perform well in real-world scenarios.
Summary: In the context of data values for AI, correctness refers to how accurately the recorded data represents the ground truth. They involve calculations related to error ratios, precision criteria, contextual error tolerance, and comparisons between dataset attribute values and their real-world values.
@@ 3.1.12. Timeliness
The timeliness of data refers to the time data collection and its relevance to the phenomenon or domain being studied in an AI application. It measures how closely the data captures the most relevant information available at the time of analysis or model training, ensuring that the data is up-to-date and reflects the present conditions. The metric may vary depending on the question being solved by an AI application. In some applications, the latest data may be needed. In others, data relevant to the pattern an AI application is trying to predict. Existing metrics define timeliness based on the existence of the most recent data.
@@
Example: Consider an AI system that predicts product demand for an e-commerce company. Timeliness of the data used for training the model would involve using the most recent sales data, customer preferences, and market trends. If the dataset contains sales data from several months ago, it may not accurately capture the current demand patterns and consumer behavior related to the current and near-future conditions. By ensuring timely data, such as incorporating daily or weekly sales updates, the AI model can better adapt to the changing market dynamics and provide more accurate demand predictions.
@@
Metrics in Literature: Two studies, one by Kaiser et al. (1970) and the other by Heinrich and Klier (2015), introduce metrics for evaluating the timeliness of attribute values. Both use probability-based approaches to assess the freshness and relevance of data. Kaiser et al.’s metric uses an exponential distribution model to calculate attribute decline rates. It indicates the average proportion of outdated attribute values within a specified time frame. It quantifies attribute age based on data quality assessment time and data acquisition time to offer an automated and interpretable measure of timeliness. In contrast, Heinrich et al. propose a probability-based currency metric (PBCM) that assesses data item timeliness using a set of probabilities. These probabilities are derived from diverse data sources and methods, including expert assessments, historical data analysis, and machine learning algorithms. While both metrics share a foundation in probability theory, Kaiser et al.’s metric focuses on attribute-level timeliness. In contrast, Heinrich et al.’s PBCM assesses data item timeliness, offering flexibility for various data types and contexts.
@@
Blake and Mangiameli (2011) propose a method to assess data timeliness using a classification model. They introduce a metric called $T$, which measures the impact of introducing new and more current data into the dataset. To evaluate data volatility and timeliness, the authors replace a percentage of old instances with new ones in the training data while assuming a fixed currency. The $T$ metric is computed based on the total quantity of records in the training and test data and the number of replacement records introduced for reclassification.
Impact on AI: Timely data is critical to AI applications trying to understand patterns, although the ‘time’ refers to the question an AI application is trying to answer. Applications trying to understand and predict patterns related to the latest trends require the most recent data.
@@
Summary: The timeliness metric assesses the recency and relevance of the data about the current state of the phenomenon or domain it represents. Kaiser et al. and Heinrich et al. introduced metrics for assessing data timeliness. Blake et al. evaluate data timeliness through a classification model by assessing data replacement’s impact on model performance.
@@ 3.1.13. Privacy Leakage
Data privacy in the context of AI refers to protecting and preserving sensitive information contained within datasets, particularly concerning the risk of unauthorized disclosure or inference of private details.
@@
A notable technique used to assess privacy is Membership Inference Attacks (MIA). MIA determines whether a specific data record was included in the training dataset used to build an AI model. By exploiting patterns and characteristics of the model’s outputs, one can infer whether or not a particular data point was part of the training set. This raises concerns about the privacy of individual data records and the potential for unauthorized access to such information.
Another dimension of AI-related privacy issues emerges with the use of synthetic data. Synthetic data is generated to mimic real data while preserving privacy by avoiding the use of actual sensitive information. However, suppose the synthetic data is too close to the real data. In that case, it can reveal private details, making it vulnerable to privacy attacks. The balance between creating useful synthetic data and ensuring privacy remains a significant challenge in AI. Evaluating the closeness of the synthetic data to the real data can be useful in determining the potential privacy leaks that can emerge during AI applications.
@@
Example: Consider a healthcare dataset used to train a machine-learning model for disease diagnosis. The dataset contains sensitive medical information about patients, including their symptoms, test results, and diagnoses. Without proper setting of privacy for data, one may determine whether a specific patient’s data was used during training, which poses a privacy risk as it could reveal a patient’s medical condition or other confidential information.
@@
Metrics in Literature: Vatsalan et al. (2022) focus on mitigating re-identification risks in released datasets within the education sector. In contrast to existing approaches that often assume prior knowledge, the proposed method employs a Markov Model to quantify re-identification risks by using all available information in the datasets, including event-level details associating multiple records with the same individual and exploring correlations between attributes.
@@
In a broader context of privacy metrics in literature, works such as SHAP R introduced by Duddu et al. (2022) and Song et al.’s (Song and Mittal, 2021) privacy risk metric are noteworthy. SHAP R quantifies the susceptibility of individual training data records to membership inference attacks by calculating Shapley values, emphasizing the influence of specific data points on model predictions. In contrast, Song et al.’s metric assesses the likelihood of a data record being present in a model’s training dataset, focusing on evaluating the privacy risk from an adversarial perspective. Both metrics aim to address privacy concerns by assessing the privacy risks associated with data records. However, they depend on the specific AI model used in the context.
@@
In another study, Carlini et al. (2022) introduced the Attack Success Rate (ASR) as a metric for privacy leakage. ASR measures the success of an attack in predicting if a specific example is part of the training dataset. It is calculated by training a model, performing an attack, and evaluating the attack’s success in correctly predicting membership. However, ASR can only be measured after training a model and conducting an attack. Alongside these contributions, Bezzi (2007), Longpré et al. (2017), and Arca and Hewett (2020) proposed entropy-based metrics to offer valuable insights into dataset anonymity to measure unpredictability and disorder in released data.
Regarding synthetic data privacy leaks, Aindo AI’s (ain, nd) privacy score assesses the privacy risk of synthetic data by comparing proximity ratios between real and synthetic datasets. The process involves calculating the Train to Train Proximity Ratio (TTPR) for real data and the Train to Synthetic Proximity Ratio (TSPR) for synthetic data. The score is derived from the ratio of records below a specific threshold in both distributions. A score of $100$ indicates minimal privacy risk, while lower scores reflect higher risk.
Impact on AI: Evaluating privacy in data impacts AI by ensuring that personal information is protected while enabling AI systems to function effectively. Privacy assessments help identify and mitigate risks associated with data breaches, intentional or unintentional or malicious accesses, and misuse, which are critical as AI technologies increasingly process personal data. Additionally, addressing privacy concerns helps avoid ethical and legal repercussions, such as biases in AI models and non-compliance with regulations like GDPR (European Parliament and Council of the European Union, 2016). Privacy evaluations also encourage the adoption of privacy design principles in data, ensuring that AI systems are developed with privacy considerations. This will ultimately lead to more ethical and responsible AI deployment.
@@
Summary: The privacy leakage metric in AI aims to safeguard sensitive information from unauthorized disclosure. Privacy metrics range from a Markov Model to address re-identification risks to SHAP R to quantify membership privacy risk using Shapley values. Attack Success Rate (ASR) is a measure of privacy leakage in membership inference attacks. Entropy-based metrics are also explored to assess dataset anonymity.
@@ 3.1.14. Sample Size
Sample size refers to the number of data points or instances selected from a population to be included in a dataset for training an AI model. It represents the subset of data used to make inferences or predictions.
@@
Example: If researchers collected data from $500$ patients to predict the likelihood of a disease based on patient characteristics, such as their age, gender, medical history, and test results, the sample size of the dataset would be $500$.
@@
Metrics in Literature: Alwosheel et al. (2018) investigate the sample size requirements for accurate decision-making analysis using Artificial Neural Networks (ANNs). They introduce a new guideline, “factor $50$,” recommending that the optimal dataset size for an ANN should be the number of adjustable parameters in the model multiplied by $50$. This guideline is more conservative than the commonly used “factor $10$” rule-of-thumb found in the literature (Haykin, 2009). However, determining the appropriate sample size for ANN-based decision-making is complex and depends on the model’s complexity, which is difficult to predict. To address this, the authors propose three approaches: evaluating ex-post if the training sample size was sufficient, using prior studies or literature to estimate the optimal number of neurons, or referring to existing literature to calculate the expected number of neurons required for the analysis.
Impact on AI: An adequate sample size ensures that AI models can learn effectively from the data, capturing the underlying patterns without overfitting or underfitting. Small sample sizes may lead to overfitting, where the model performs well on training data but poorly on new, unseen data due to learning noise or random variations rather than the true signal. Conversely, excessively large sample sizes can be resource-intensive and may not significantly improve model performance beyond a certain point. Rajput et al. (2023) have shown varying impacts of sample size on model accuracy, with some models performing better with larger datasets, while others may not see significant gains past a certain threshold.
@@
Summary: In AI, sample size denotes the quantity of data points chosen from a population for analysis or model training. The “factor $50$” determines sample size in decision-making with Artificial Neural Networks (ANNs), contrasting with the commonly used “factor $10$”.
3.1.15. FAIR Principle Compliance Score